Find 3 consecutive numbers where the product of the smaller two numbers is 19 less than the square of the largest number.
Answer by JoelSchwartz(130)

x=first consecutive number
y=second consecutive number
z=third consecutive number
y=x+1
z=x+2
xy=z^2-19
x(x+1)=(x+2)(x+2)-19
x^2+x=x^2+4x+4-19
15=3x
x=5
y=6
z=7